OWL Question for Quiz #20: Module 10 Problems 140 PM Question Status AIA Inc is looking to manage its cash position using the EO0 model The company is consuming cash at the rate of $5500 per day, and is open for days in s 365 the year ach tme the rm sells secur testo obtain the cash, t costs them S150 The interest rate is 2.10% What are the total costs (and al storage with the EOQ? CHECK ANSWER !
Answer #1

Please provide rating..

EOQ = square root of (2* annual consumption*transaction cost/Interest rate)
Annual consumption =5500*365 2007500
Transaction cost 150
Interest rate 2.10%
EOQ = =square root of (2*2007500*150/2.1%)
EOQ = 169347
Average cash balance = 169347/2       84,673.74
Therefore storage cost = 84673.74*2.1%         1,778.15
Number of transaction = 2007500/169347              11.85
Transaction cost = 11.85*150         1,778.15
Total cost = 1778.15+1778.15         3,556.30